# MeshCone: Second-Order Cone Programming for Geometrically-Constrained Mesh Enhancement

### v3

@@ Abstract
Modern geometric generation methods rely heavily on deep learning methods that, while powerful, often lack interpretability and require extensive training data. This work introduces MeshCone, a convex optimization framework for mesh enhancement from partially deformed meshes that requires no training data. We formulate the problem as a second-order cone program where vertex positions are optimized to align with target geometry while enforcing smoothness through convex edge-length regularization. Our convex relaxation enables deterministic, interpretable solutions with proven convergence properties via the Splitting Conic Solver (SCS). We demonstrate robust performance across 56 diverse object categories from ShapeNet and ThreeDScans, achieving superior refinement quality compared to classical baselines while maintaining sub-second inference times. This work establishes a principled baseline demonstrating what convex optimization alone can achieve, providing mathematical guarantees and interpretability that complement data-driven approaches.
@@ 1 Introduction
The field of computer graphics has been deeply engaged with mesh generation since the introduction of the Marching Cubes algorithm [10.1145/37402.37422] nearly four decades ago. Over the years, diverse techniques have emerged—from classical methods based on Signed Distance Fields (SDFs) and Radial Basis Functions (RBFs) to modern learning-based approaches using neural networks and autoregressive architectures. Despite this rich history, a fundamental question remains largely unexplored: What refinement quality can be achieved using pure convex optimization with mathematical guarantees?
Modern mesh generation increasingly relies on data-driven methods that learn geometric priors from large datasets. While these approaches achieve impressive results, they introduce challenges in interpretability, require extensive training data, and lack formal guarantees about solution quality or convergence behavior. This raises a natural question: what can mathematical optimization alone accomplish? Understanding this baseline is essential for evaluating whether the added complexity of learning-based methods is justified and for identifying opportunities where hybrid optimization-learning approaches might excel.
Convex optimization has long been a cornerstone of fields such as finance, engineering, and economics, offering a principled framework that guarantees globally optimal solutions under well-defined constraints. Its efficiency, robustness, and optimality make it desired for applications where stability and reproducibility are essential. However, within computer graphics, the potential of convex optimization for direct mesh generation remains largely unexplored. While optimization techniques are common in mesh processing tasks—such as parameterization, deformation, and smoothing—few works investigate whether convex formulations can serve as viable standalone approaches for mesh optimization.
This work establishes what is achievable through pure convex optimization for mesh refinement from partially deformed or noisy input meshes that can be generated from any mesh generation approach that produces suboptimal representations. To address these quality issues, we introduce MeshCone, a conic optimization framework that formulates mesh refinement as a second-order cone program (SOCP). Our approach operates directly on vertex positions, enforcing geometric consistency through convex edge-length constraints and achieving smoothness via Tikhonov regularization. By employing the Splitting Conic Solver (SCS), we obtain solutions with provable global optimality, guaranteed convergence at rate $O(1/k)$, and Lipschitz stability under perturbations—properties unavailable in heuristic or neural approaches.
Unlike learning-based pipelines, this formulation requires no training data, introduces no learned parameters, and provides deterministic, interpretable solutions where every constraint has clear geometric meaning. We demonstrate that it provides a principled, mathematically grounded baseline with theoretical guarantees of global optimality and predictable convergence.
@@
  • Baseline establishment: We provide a deterministic, training-free baseline with mathematical guarantees against which future optimization-based and hybrid methods can be compared, identifying both capabilities and limitations of pure convex approaches.
@@ 4 Experiments
The approach is specifically evaluated on 53 meshes from the ShapeNet dataset [chang2015shapenetinformationrich3dmodel], with each mesh representing a distinct object category. This diverse evaluation set spans the full range of ShapeNet taxonomy, including vehicles, furniture, electronics, kitchenware, and many other categories. This setup allows us to assess whether the optimization process can generalize across vastly different structures. In addition, we tested our solution against Laplacian smoothing on three meshes generated from the work of [Binninger_2025], which are highly detailed and contain thousands of faces and vertices, making them ideal for a real-world test [Laric2012ThreeDScans]. We sampled 10000 points from these representations to obtain the respective centroids used for alignment between the deformed vertices and the ground truth mesh.
@@ 6.1 Limitations
While the convex optimization framework provides strong theoretical guarantees, it occasionally struggles with preserving extremely high-frequency details in complex geometric regions. These limitations could be addressed in future work through specialized constraint formulations that explicitly target fine-scale features.
@@ 8 Conclusions
This work introduced a convex optimization framework for mesh improvement that relies exclusively on linear constraints and mathematically interpretable solvers. Through this research, we demonstrated that high-quality developments can be achieved while preserving the intrinsic geometric and topological properties of the input meshes. Our tests on ShapeNet and ThreeDScans confirmed that the proposal really enhances meshes by improving the connectivity among the vertices and edges while including smoothness into the final representation. MeshCone enables user to refine meshes without requiring complex hyperparameters or long training times, what makes it suitable for different applications related to geometry processing and meshing.
We believe that the future of geometric deep learning lies not in replacing optimization with learning, but in principled integration of both paradigms. This work provides a rigorous baseline and mathematical foundation upon which such hybrid methods can be built, offering a path toward generative models that combine creativity with geometric correctness.

### v4

@@ Abstract
Modern mesh generation pipelines—whether learning-based or classical—often produce outputs requiring post-processing to achieve production-quality geometry. This work introduces MeshCone, a convex optimization framework for guided mesh refinement that leverages reference geometry to correct deformed or degraded meshes. We formulate the problem as a second-order cone program where vertex positions are optimized to align with target geometry while enforcing smoothness through convex edge-length regularization. MeshCone performs geometry-aware optimization that preserves fine details while correcting structural defects. We demonstrate robust performance across 56 diverse object categories from ShapeNet and ThreeDScans, achieving superior refinement quality compared to Laplacian smoothing and unoptimized baselines while maintaining sub-second inference times. MeshCone is particularly suited for applications where reference geometry is available, such as mesh-from-template workflows, scan-to-CAD alignment, and quality assurance in asset production pipelines.
@@ 1 Introduction
The field of computer graphics has been deeply engaged with mesh generation since the introduction of the Marching Cubes algorithm [10.1145/37402.37422] nearly four decades ago. Over the years, diverse techniques have emerged—from classical methods based on Signed Distance Fields (SDFs) and Radial Basis Functions (RBFs) to modern learning-based approaches using neural networks and autoregressive architectures. Yet regardless of the generation method, a persistent challenge remains: how can we efficiently refine imperfect meshes toward high-quality target geometry with mathematical guarantees?
Modern mesh generation pipelines—whether classical or learning-based—frequently produce outputs that require post-processing before deployment. Generated meshes may exhibit noise, geometric distortions, or deviation from intended shapes. While learning-based refinement methods can address these issues, they require extensive training data, lack interpretability, and provide no formal guarantees about solution quality or convergence. This motivates the need for principled refinement approaches that can leverage available reference geometry to correct mesh defects reliably and predictably.
Convex optimization has long been a cornerstone of fields such as finance, engineering, and economics, offering a principled framework that guarantees globally optimal solutions under well-defined constraints. Its efficiency, robustness, and optimality make it desirable for applications where stability and reproducibility are essential. Within computer graphics, optimization techniques are common in mesh processing tasks—such as parameterization, deformation, and smoothing—yet few works investigate whether convex formulations can serve as effective guided refinement methods that leverage target geometry to achieve provably optimal corrections.
This work introduces MeshCone, a conic optimization framework for geometry-guided mesh refinement. Given a source mesh and reference target geometry, MeshCone formulates refinement as a second-order cone program (SOCP) that optimizes vertex positions to align with the target while enforcing smoothness via convex edge-length constraints and Tikhonov regularization. By employing the Splitting Conic Solver (SCS), we obtain solutions with provable global optimality, guaranteed convergence at rate $O(1/k)$, and Lipschitz stability under perturbations—properties unavailable in heuristic or neural approaches.
Unlike blind smoothing methods such as Laplacian filtering that uniformly regularize without geometric awareness, MeshCone performs target-aware optimization that preserves fine details present in the reference while correcting structural defects in the source mesh. This makes our approach particularly suited for workflows where reference geometry is available, including:
@@
  • Baseline establishment: We provide a deterministic, geometry-aware baseline with mathematical guarantees against which future optimization-based and hybrid methods can be compared, identifying both capabilities and limitations of pure convex approaches.
@@ 4 Experiments
The approach is specifically evaluated on 53 meshes from the ShapeNet dataset [chang2015shapenetinformationrich3dmodel], with each mesh representing a distinct object category. This diverse evaluation set spans the full range of ShapeNet taxonomy, including vehicles, furniture, electronics, kitchenware, and many other categories. This setup allows us to assess whether the optimization process can generalize across vastly different structures. In addition, we tested our solution against the classical Laplacian smoothing on three meshes generated from the work of [Binninger_2025], which are highly detailed and contain thousands of faces and vertices, making them ideal for a real-world test [Laric2012ThreeDScans]. We sampled 10000 points from these representations to obtain the respective centroids used for alignment between the deformed vertices and the ground truth mesh.
@@ 6.1 Limitations
While the convex optimization framework provides strong theoretical guarantees, it occasionally struggles with preserving extremely high-frequency details in complex geometric regions. The global nature of our centroid-alignment objective and edge-length constraints can over-smooth localized details. These limitations could be addressed in future work through localized optimization that decomposes the mesh into regions with independent convex subproblems, or multi-scale refinement strategies that apply hierarchical optimization at different resolution levels.
@@ 8 Conclusions
This work introduced MeshCone, a convex optimization framework for geometry-guided mesh refinement that leverages reference geometry to achieve provably optimal corrections. Through this research, we demonstrated that high-quality refinement can be achieved while preserving the intrinsic geometric and topological properties of the input meshes. Our tests on ShapeNet and ThreeDScans confirmed that the framework enhances meshes by improving vertex alignment and edge consistency while enforcing smoothness in the final representation. MeshCone enables users to refine meshes with minimal hyperparameter tuning and no training requirements, making it suitable for workflows where reference geometry is available, such as template-based modeling, scan-to-CAD alignment, and asset production pipelines.
We believe that the future of geometric deep learning lies not in replacing optimization with learning, but in principled integration of both paradigms. This work provides a rigorous mathematical foundation upon which hybrid methods can be built, offering a path toward systems that combine the expressiveness of generative models with the guarantees of convex optimization.